Find the solution of the equation x3-2x2-3x in z12

1
Expert's answer
2021-05-31T17:41:02-0400

First let us factorize this polynomial :

"x^3-2x^2-3x=x\cdot(x-3)\cdot(x+1)"

The obvious solutions are "0; 3; 11" (as "11=-1" in "\mathbb{Z}_{12}" ). However, as "\mathbb{Z}_{12}" contains the zero divisors, there could possibly be other solutions.

"12=4\cdot 3" and thus if "x" is a solution of this equation, we should have "3|x" or "3|(x+1)" (as the condition "3|(x-3)" is equivalent to "3|x"). Therefore, the other possible solutions may be "6;9;2;5;8"

Now we can also see, that if "x" is even, then "x-3, x+1" are odd and thus if "x" is a solution and it is even, it should be divisible by 4. However, if "x" is odd, then both "x-3, x+1" are even, so their product is divisible by 4. These conditions yield that the solution set is

"S:= \{0;3;5;8;9;11 \}"